The equation 3x − 2y = −16 can be used to determine the size of a recommended refrigerator, where y represents the volume of the refrigerator in cubic feet, and x represents the number of people (greater than 2) living in the residence. When this equation is graphed on a coordinate plane, where does the line intercept the y-axis?
Answer by stanbon(57387)

The equation 3x − 2y = −16 can be used to determine the size of a recommended refrigerator, where y represents the volume of the refrigerator in cubic feet, and x represents the number of people (greater than 2) living in the residence. When this equation is graphed on a coordinate plane, where does the line intercept the y-axis?
---
The y-axis passes thru all the points where x = 0.
So, let x = 0 and solve your equation for "y":
---
3x-2y = -16
-2y = -16
y = 8
---
The y-intercept is (0,8)
===========================
Cheers,
Stan H.
===========